OpenStudy (anonymous):

f(x)=3-absolutevalue(x-3) find the derivative.

OpenStudy (anonymous):

Rewrite it as a piece-wise function\[ f(x) = 2-|x-3| = \begin{cases} 2-(x-3) &=&-x-1 &x>0\\ 2- (-(x-3))& =&x-5 &x<0 \end{cases}\] Now we can differentiate each piece. \[ f'(x) = \begin{cases} (-x-1)' &=&-1 &x>0\\ (x-5)'&=&1 &x<0 \end{cases} \]

OpenStudy (anonymous):

Whoops it should say: \[ f'(x) = \begin{cases} (-x-1)' &=&-1 &x>3\\ (x-5)'&=&1 &x<3 \end{cases} \]

OpenStudy (anonymous):

\[ f(x) = 2-|x-3| = \begin{cases} 2-(x-3) &=&-x-1 &x-3>0 \implies x>3\\ 2- (-(x-3))& =&x-5 &x-3<0 \implies x<3 \end{cases} \]

OpenStudy (anonymous):

I'm trying to determine if I can use Rolle's Theorem for this function... I'm still unsure. Oh and this is on the interval [0,6]

OpenStudy (anonymous):

What are you trying to show?

OpenStudy (anonymous):

My instructions are: Determine whether Rolle's Theorem can be applied to f on the closed interval [a,b]. If Rolle's Theorem can be applied, find all values of c in the open interval (a,b) such that f'(c)=0.

OpenStudy (anonymous):

The problem with Rolle's theorem is that it requires the function be differentiable on the interval, which in this case isn't true. \(f(x)\) is not differentiable at 3.

OpenStudy (anonymous):

Basically your interval can't contain \(3\).

OpenStudy (anonymous):

Okay, I sort of get it...

OpenStudy (anonymous):

Anyway it's pretty obvious that \(f'(x)\neq 0\) for all \(x\) since we know it is either \(-1\) or \(1\).